The sides of a triangle are in a ratio of 4:5:6. Solve for the smallest angle


1
Expert's answer
2021-10-18T16:11:54-0400

Let the angles be ;4x,5x,6x

Sum of interior angles in a triangle, adds up to 180 °.

Thus:"4x+5x+6x=180"

"15x=180"

"x=12"

Hence the smallest angle =12×4 =48°